# BIFROST: A First-Principles Model of Polarization Mode Dispersion in Optical Fiber

### v1

@@ VIII Example Simulation: Wavelength-Division Multiplexing for PMD Compensation in Quantum Networks
Several PMD compensation schemes have been investigated and demonstrated. One class of compensation schemes, which has its roots in decades of work on networks dedicated to quantum key distribution (QKD), uses resources available in QKD (such as measuring the qubit error rate or using the discarded qubits) to perform compensation [25, 26, 27]. In more flexible quantum networks, which do not make any assumption about the kinds of tasks performed by the network [22, 105], compensation schemes generally fall into two categories, as shown in Fig. 7. One type is time-division multiplexing (TDM). In TDM, the quantum communication channel is paused long enough for classical reference pulses with known polarizations to be sent through the fiber. The polarization of these pulses is measured at the fiber output, and that information is used to correct the polarization rotations experienced by the quantum information. In between reference measurements, the network has no knowledge of the PMD drift, and the reference measurements and drift compensation must be updated regularly. TDM is easy to implement , and has been used for decades [e.g. 28, 29, 30, 31, 32, 33, 34, 35, 36]. However, it comes with the costs of network downtime and degraded fidelity in between measurements. (Ref. [10], mentioned above, used TDM and demonstrated a network uptime of 99.83% with average entanglement distribution fidelity of , indicating that, for “quiet” fibers, TDM may produce good performance.)
@@
This example simulation only scratches the surface of possible WDM compensation investigation, but it suggests that there are regimes in which WDM schemes may work well. It will be important to formulate more comprehensive metrics for compensation fidelity, investigate other sources of environmental variation, explore how the results change with the total PMD of the fiber, simulate extensions to the basic WDM scheme presented here [39, 40], and compare WDM to TDM more directly to determine whether WDM is advantageous.

### v2

@@ VII BIFROST Usage and Recommended Operating Limits
As a final usage note for BIFROST, modeling fibers so carefully makes BIFROST reminiscent of digital twin models [107, 116] for optical fiber networks. While digital twins are well-established for wireless communication [e.g. 58], it remains challenging to develop practical digital twins for fiber-based networks, and research in this area is ongoing [e.g. 39, 122, 17]. For tracking state of polarization specifically, a digital twin is difficult to implement because the SOP at the end of the fiber is the accumulation of time-varying rotations along a fiber that result from several birefringence mechanisms. Parameters that could be used to predict the SOP would be challenging to acquire in deployed fiber. Thus digital twins are generally impractical for our purposes. Rather, BIFROST provides useful physics-based information about the statistical performance of a fiber.
@@ VIII Example Simulation: Wavelength-Division Multiplexing for PMD Compensation in Quantum Networks
Several PMD compensation schemes have been investigated and demonstrated. One class of compensation schemes, which has its roots in decades of work on networks dedicated to quantum key distribution (QKD), uses resources available in QKD (such as measuring the qubit error rate or using the discarded qubits) to perform compensation [35, 98, 67]. In more flexible quantum networks, which do not make any assumption about the kinds of tasks performed by the network [7, 109], compensation schemes generally fall into two categories, as shown in Fig. 7. One type is time-division multiplexing (TDM). In TDM, the quantum communication channel is paused long enough for classical reference pulses with known polarizations to be sent through the fiber. The polarization of these pulses is measured at the fiber output, and that information is used to correct the polarization rotations experienced by the quantum information. In between reference measurements, the network has no knowledge of the PMD drift, and the reference measurements and drift compensation must be updated regularly. TDM is easy to implement , and has been used for decades [e.g. 72, 71, 85, 119, 110, 1, 101, 28, 108]. However, it comes with the costs of network downtime and degraded fidelity in between measurements. (Ref. [30], mentioned above, used TDM and demonstrated a network uptime of 99.83% with average entanglement distribution fidelity of , indicating that, for “quiet” fibers, TDM may produce good performance.)
@@
This example simulation only scratches the surface of possible WDM compensation investigation, but it suggests that there are regimes in which WDM schemes may work well. It will be important to formulate more comprehensive metrics for compensation fidelity, investigate other sources of environmental variation, explore how the results change with the total PMD of the fiber, and compare WDM to TDM more directly to determine whether WDM is advantageous. Additionally, BIFROST may be useful in simulating more creative variations on WDM. For instance, Refs. [38, 114] proposed using two reference wavelengths on opposite sides of the signal wavelength, and Ref. [27] used heterodyne detection to relax power requirements on the WDM reference, resulting in negligible added noise from nonlinear scattering.